OpenStudy (anonymous):

Find the equation of the line that has a slope of 5 and passes through the point (-8,6).

OpenStudy (anonymous):

This math is so pointless.

OpenStudy (johnweldon1993):

Point slope form... \[\large y - y_1 = m(x - x_1)\] Your \(\large m = 5\) Your \(\large y_1 = 6\) Your \(\large x_1 = -8\) so \[\large y - 6 = 5(x + 8)\] Distribute the 5 into the parenthesis... \[\large y - 6 = 5x + 40\] Add 6 to both sides... \[\large y = 5x + 46\]

OpenStudy (anonymous):

You.. are so stupid. I meant that it's stupid, and ridiculous, and there's NO POINT IN DOING IT.!